1.The siren of a firetruck has a frequency of 1.20 kHz. When the firetruck is moving at 24.6 m/s east, what frequency is heard by someone in a car traveling 4.47 m/s east away from the fire truck. The temperature is 30.0 C

2. A container holds 0.320 m3 of oxygen at an absolute pressure of 6.00 atm. A valve is opened allowing 1/3 of the gas to escape. If the volume of the container and the temperature remain constant throughout the process, what is the final pressure of the gas?
Answer #1

o Vj = (331 331 +0.61x30 ) = 349.3 mg 349.3 + 24.6 of -) (1200) = [1268.28 Hz 349.3+ 4.47 2 pr=nRT Pan Pi Pa n₂ = 2h 3 ni n